# Overcoming the Zero-Rate Hashing Bound with Holographic Quantum Error Correction

### v1

@@ Abstract
Several recent techniques for modifying topological codes with single-qubit Clifford operators have shown high resilience against pure Pauli noise. Paramount to these findings has been the demonstration that several variants exhibit error thresholds often attaining or exceeding the zero-rate hashing bound, a known benchmark for code-capacity noise channels, for biased noise. Additionally, direct comparison with the hashing bound has shown that several topological codes outperform the hashing bound at points of finite Pauli noise biases. Motivated by these observations, we study zero-rate holographic quantum error correction codes, discovering very high threshold values under diverse and finitely-biased noise channels using a tensor-network decoding approach. Our results establish that all codes tested achieve or surpass the hashing bound at various points, ranging from pure 2-Pauli noise ($\eta=0$) to pure 1-Pauli noise ($\eta=+\infty$), thereby demonstrating that holographic codes exhibit excellent error tolerance in the code-capacity picture. Such findings imply the existence of a structured and systematic method for constructing high-threshold codes suitable for realistically motivated noise channels. To our knowledge, this work is also the first instance of such remarkable threshold behavior in stabilizer quantum codes for the pure 2-Pauli noise regime, as well as for finitely-biased noise channels.
@@ I Introduction
Utilizing principles from quantum error correction (QEC) is thus far the most viable path towards the construction of a large-scale, fault-tolerant quantum computer [1, 2, 3]. In the field of QEC, stabilizer codes [4] as well as topological quantum codes [5] have come to dominate the domain, of which ever more complex instances are being realized in experimental demonstrations [6, 7, 8]. In order to inform these experiments, simulations of error thresholds for entire classes of stabilizer codes have been implemented; here, the family of surface codes has become particularly prominent, due in large part to robust error thresholds under various error models of progressively more realistic circumstances, as well as their known practical appeal from local stabilizer checks and planar architectural layout [9, 5]. On the one hand, it is known that simple noise channels such as the quantum erasure and depolarizing channels do give preliminary information on how resilient a quantum code may be at the theoretical level [10, 11]. However, in practice, the experimental calculus is much different, with noise often biased towards a certain subset of Pauli errors, such as finitely-biased or pure $Z$ noise found in qubit technologies which experience dephasing noise [12, 13, 6, 14, 15, 16], or a combination of several Pauli errors at once, as in the square-lattice Gottesman-Kitaev-Preskill (GKP) code [17, 18, 19, 20]. It is in this area that topological codes (e.g. Clifford-deformed topological codes such as the XZZX surface code), concatenated codes based on repetition codes, as well as graph states yield exceptionally high thresholds for pure 1-Pauli $X$, $Y$, or $Z$ noise, achieving $50\%$ in many cases [21, 22, 23, 24, 25, 26, 27, 28]. In these studies, the assumptions necessary for biased-noise code-capacity explorations traditionally include independent and identically distributed (IID) single-qubit Pauli channels
$\mathcal{E}(\rho)=(1-p)\rho+p(r_{X}X\rho X+r_{Y}Y\rho Y+r_{Z}Z\rho Z)\ ,$ (1)
with the relative error probabilities $r_{X}+r_{Y}+r_{Z}=1$, typically in the regime of $r_{X}=r_{Y}$ with a bias $\eta=\frac{r_{Z}}{r_{X}+r_{Y}}\geq\frac{1}{2}$ (in the case of $Z$-biased noise). When $\eta=\frac{1}{2}$ one recovers the standard depolarizing noise channel. Additionally, the behavior of these $Z$-biased thresholds was studied, and it was found that such surface-code variants in many cases either match or exceed the error threshold bounds predicted by the zero-rate hashing bound [11].
@@
In this work, we demonstrate that such intriguing behavior is also present in holographic quantum codes [29, 33, 32, 30, 31]. Using the tensor-network decoder described in [31, 34], we study the biased-noise error thresholds in detail for several zero-rate holographic codes. Such codes are derived from elemental seed tensors describing a simple code, contracted in a tensor network with hyperbolic geometry (Fig. 1). Here we test the Harlow-Preskill-Pastawski-Yoshida (HaPPY) model [29], whose seed tensor corresponds to the $\llbracket 5,1,3\rrbracket$ code. We additionally test: the Steane code [32]; the $\llbracket 6,1,3\rrbracket$ code [31]; and the holographic surface-code fragment (SCF) [30]. For all of the codes tested, we find evidence of highly resilient code-capacity thresholds and display characteristics of their resulting ternary biased-noise plots. Key results in our work include: 1) Evidence that the zero-rate HaPPY code exhibits $50\%$ error thresholds for pure biased Pauli noise. 2) The holographic Steane code obtains a depolarizing noise threshold of $18.985\%$, competitive with known code-capacity bounds derived for several topological codes using a mapping to statistical-mechanical models [35]; 3) For several pure 2-Pauli noise channels (i.e., consisting of a subset of two out of the three Pauli errors), the zero-rate SCF and $\llbracket 6,1,3\rrbracket$ holographic codes exceed current state-of-the-art results [28, 36], in addition to surpassing the hashing bound. 4) Finally, for all of the codes evaluated, error thresholds were found to overtake the bound for numerous finitely-biased points on the hashing curve, within the bias range of $0\leq\eta\leq+\infty$. Our results strongly suggest that holographic quantum codes possess high resilience against various types of noise, bolstering the case that these codes may find utility in practical quantum information processing.
II Holographic Quantum Error Correction
[…]
@@ II Holographic Quantum Error Correction
Originally developed to model the AdS/CFT correspondence [45, 46, 47, 48, 49], holographic quantum error correction involves encoding maps from bulk (logical) degrees of freedom on a $d{+}1$-dimensional hyperbolic space to $d$-dimensional boundary (physical) degrees of freedom. While in full AdS/CFT the bulk and boundary degrees of freedom are associated with weakly-coupled quantum gravity on an asymptotically anti-de Sitter (AdS) background and strongly-coupled conformal field theory (CFT), respectively, aspects of this duality for $d=1$ can be captured by simple tensor network codes [50, 29, 51, 33]. From a practical perspective, there are many reasons to find the concept of a holographic quantum code appealing. Firstly, holographic codes are relatively simple to construct and scale from small, atomized examples to larger example codes. As a direct consequence of the quantum LEGO formalism [52], concepts such as code structure and transversal logical operations are straightforward to intuit. Secondly, the boundary of holographic codes exhibits a quasiperiodic self-similarity [53, 54], allowing one to rescale the number of physical and logical qubits using local inflation rules that also determine the code’s rate and distance scaling [34, 55]. Thirdly, encoding rates of holographic codes are known to be quite high and tunable [34], competitive with many recent proposals of quantum low-density parity-check (qLDPC) codes [56]. Fourthly, it is known that most seed tensors for holographic codes can be easily mapped to graph states, implying that efficient preparation schemes likely exist [57, 58, 59, 60]. Lastly, and perhaps most consequentially, it has been shown that holographic codes exhibit high resilience against various noise channels. Indeed, it was first shown in [29] that holographic quantum error correction codes exhibit high thresholds against the quantum erasure channel; subsequently, several works have shown the capability of certain codes constructions to protect against depolarizing and 1-Pauli noise, potentially as well as topological codes [34, 32, 30, 31, 61, 55]. However, to our knowledge, a systematic biased-noise threshold study for even zero-rate holographic codes has not been performed, let alone for their constant-rate versions111 [62] discusses biased noise in finite-rate HaPPY pentagon codes, but it only captures the non-detectable error probability for a code of fixed length.. As such, our work takes the first step towards understanding holographic codes under more generalized noise channels.
@@
The main ingredient of a holographic quantum code is the seed tensor defining the encoding map for a single tile. Here we consider the zero-rate case with a logical qubit only on the central tile; on the remaining tiles all tensor legs are planar, effectively dispersing the logical information towards the tiling boundary. Note that in generic holographic codes, each tile may hold a logical qubit, leading to a nonzero asymptotic rate as the number $L$ of layers is increased. While copies of the same seed tensor are typically used in this case, one may also employ several different seed tensors [63] in a spirit similar to heterogeneous concatenation methods for tree-style concatenated codes [64, 65]. In Appendix A, we provide a table with all of the relevant information about each seed tensor for each code utilized in this work.
@@ III Results
Threshold calculations were performed using the tensor-network decoder from [31, 34]. These decoders and others will be made public in an upcoming software package for holographic quantum error correction codes [66]. However, for the purposes of understanding the current work, we focus on surveying biased-noise resilience for a swath of known holographic codes, and not on a systematic treatment of the tensor-network decoding methodology [67].

### v2

@@ Abstract
A crucial insight for practical quantum error correction is that different types of errors, such as single-qubit Pauli operators, typically occur with different probabilities. Finding an optimal quantum code under such biased noise is a challenging problem, related to finding the (generally unknown) maximum capacity of the corresponding noisy channel. A benchmark for this capacity is given by the hashing bound, describing the performance of random stabilizer codes, which leads to the challenge of finding codes that reach or exceed this bound while also being efficiently decodable. In this work, we show that asymptotically zero-rate holographic codes, built from hyperbolic tensor networks that model holographic bulk/boundary dualities, fulfill both conditions. Of the five holographic code models considered, all are found to reach the hashing bound in some bias regime and one, the holographic surface-code fragment, appears to even exceed the capacity of previously known codes in the 2-Pauli-dominated noise regime. In addition, we consider Clifford deformations that allow all considered codes to reach the hashing bound for 1-Pauli-dominated noise as well. Our results thus establish that holographic codes, which were previously shown to possess efficient tensor-network decoders, also exhibit competitive thresholds under biased noise.
@@ I Introduction
Understanding and mitigating the impact of quantum noise is crucial for quantum computation and communication. The most common way to characterize quantum noise is through quantum channels, which provide rigorous descriptions for how a quantum state is transformed or transmitted in the presence of errors. The maximum number of qubits that can be transmitted reliably in a noisy environment is known as the quantum channel capacity (QCC) [1, 2, 3, 4, 5, 6], which provides an important theoretical limit for quantum computation and communication tasks. However, determining the precise value of the QCC has been incredibly challenging except for a few isolated instances [7]. In particular, the QCC is not accurately known even for a channel as ubiquitous as the depolarizing channel. A large part of the difficulty is related to the potential superadditivity of coherent information, which hinders a direct efficient evaluation of the quantum capacity when the channel acts independently and identically over many qubits. As a result, finding lower and upper bounds of the quantum capacity constitute important progress in quantum Shannon theory [8, 1].
Aside from theoretical limits, it is also of practical interest to design classes of quantum error correction (QEC) codes that can come arbitrarily close to the QCC, thus maximizing the protection of information during noisy transmission or computation. The leading method for generating such practical QEC codes is to consider stabilizer codes, which are closely connected to classical linear codes [9, 10, 11]. However, unlike random classical linear codes which are typically capacity-achieving for symmetric channels, random quantum stabilizer codes do not generally achieve the optimal transmission rate for their analogous quantum counterparts such as the depolarizing channel or other Pauli channels. Instead, they attain a theoretical lower bound on QCC called the hashing bound [12, 13, 8]. Currently, there remains a large gap between the theoretical QCC upper bound and the hashing lower bound. In fact, only tiny improvements over the hashing bound have been made over the years for Pauli channels [2, 14, 15]. Furthermore, optimal decoding of random stabilizer codes is #$\#P$-complete [16], severely limiting their applicability. Therefore, it is of both theoretical and practical interest in quantum coding theory to identify efficiently decodable codes that can attain or even exceed the hashing bound. A class of such codes will not only provide new approaches for obtaining tighter QCC lower bounds but also inspire proposals for more efficient strategies of error correction.
Along this direction, a significant amount of work has been carried out over the past few decades, which has led to a number of QEC codes that approach the hashing bound, e.g. quantum turbo codes and related variations [17, 18, 19]. More recently, certain topological codes (e.g. Clifford-deformed topological codes such as the XZZX surface code) and certain concatenated codes built from various-sized repetition codes have also been able to attain or exceed the error threshold set by the hashing limit [20, 21, 22, 23, 24, 25, 26, 27].
In this work, we demonstrate that asymptotically zero-rate holographic quantum codes, a relatively recent code class derived from properties of the anti-de Sitter / conformal field theory correspondence (AdS/CFT) [28, 29, 30, 31, 32, 33, 34, 35], can achieve and supersede the hashing bound for various Pauli channels of interest. Using the tensor-network decoder described in [36, 37], we study zero-rate versions of five different holographic codes: the Harlow-Preskill-Pastawski-Yoshida (HaPPY) code [33]; the tailored $\llbracket 7,1,3\rrbracket$ code [38]; the holographic Steane code [39]; the holographic $\llbracket 6,1,3\rrbracket$ code [36]; and the holographic surface-code fragment (SCF) [40].
Our main results can be stated as follows. In the limit of pure 2-Pauli noise channels, the thresholds we have calculated for zero-rate SCF and $\llbracket 6,1,3\rrbracket$ holographic codes exceed current state-of-the-art results [27, 15, 2]. To our knowledge, no other class of quantum code has thus far been able to supersede the hashing limit for pure 2-Pauli noise. In fact, all of the holographic codes tested in this limit either come within $2\%$ of, attain, or overcome the hashing bound completely. Moreover, all of the codes tested were found to either attain or overtake the hashing bound for numerous distinct Pauli channels. Finally, we also apply the method of Clifford deformations from [21, 20] to the holographic Steane code, demonstrating that the threshold of the code can be drastically modified for depolarizing, pure X, and pure Z noise channels. Our results provide strong evidence that holographic codes possess very high resilience against various types of Pauli noise channels and can be tuned accordingly. In sharp contrast with topological codes, maximum likelihood decoding of holographic codes and the computation of weight enumerator polynomials are possible in polynomial time, even with exact tensor network contractions [36, 41], further bolstering their practical potential for both quantum error correction and for numerical studies of QCC bounds.
The remainder of this article is organized as follows: we review in the background (Section II) the basics of holographic codes and their properties for practical QEC in Section II.1, as well as an explanation for how tensor-network decoding functions (Section II.2). In Section III, we begin by discussing first the experimental setup for our simulations, the codes tested, and the biased-noise points considered (Section III.1); in Section III.2, we present all of the results of our study. Finally, we close with a discussion of future directions and implications of our work (Section IV), followed by a discussion of possible practical applications for holographic codes (Section IV.1).
II Background
II.1 Holographic Quantum Error Correction
@@
Holographic quantum codes at their core are subsystem stabilizer codes [42, 43, 44]. Stabilizer codes themselves are defined as codes for which the logical operators $\mathsf{L}$ and codeword stabilizers $\mathsf{S}$ consist of elements $P_{i}\in\mathsf{P}^{n}$, where $\mathsf{P}^{n}$ is the $n$-qubit Pauli group, $P_{i}$ takes the form $P_{i_{1}}\otimes\cdots\otimes P_{i_{n}}$, and $P\in\{I,X,Y,Z\}$, i.e., the single-qubit Pauli operators. In subsystem stabilizer codes such as holographic codes, one typically can choose to subdivide the physical Hilbert space as $\mathsf{H}_{p}=\mathsf{H}\otimes\mathsf{\bar{H}}$, where $\mathsf{H}$ represents the code subspace. It is also given that $\mathsf{H}$ can be decomposed as $\mathsf{H}=\mathsf{H}_{L}\otimes\mathsf{H}_{G}$, where $\mathsf{H}_{L}$ and $\mathsf{H}_{G}$ represent the logical and gauge subsystems of the code subspace, respectively. Holographic codes were first treated as such subsystem stabilizer codes in [45, 43, 46]. It is also wholly possible to consider finite-rate versions of holographic codes, wherein individual seed tensors can be specified with logical-qubit implantation. In this work, we consider only zero-rate versions of holographic codes, wherein only the central logical index remains (Figure 1). Additionally, we mention briefly that it is possible to construct non-Abelian stabilizer, non-stabilizer, or even approximate quantum codes in the holographic context [47, 48, 45, 49, 50]; however, we shall limit ourselves to stabilizer versions of holographic codes in the present work.
Originally developed to model the AdS/CFT correspondence [28, 29, 30, 51, 52], holographic quantum error correction involves encoding maps from bulk (logical) degrees of freedom on a $d{+}1$-dimensional hyperbolic space to $d$-dimensional boundary (physical) degrees of freedom. While in full AdS/CFT the bulk and boundary degrees of freedom are associated with weakly-coupled quantum gravity on an asymptotically anti-de Sitter (AdS) background and strongly-coupled conformal field theory (CFT), respectively, aspects of this duality for $d=1$ can be captured by simple tensor network codes [31, 33, 34, 35].
@@
From a practical perspective, there are many reasons to find the concept of a holographic quantum code appealing. Firstly, holographic codes are relatively simple to construct and scale from small, atomized examples to larger example codes. As a direct consequence of the quantum LEGO formalism [53], concepts such as code structure and transversal logical operations are straightforward to intuit. Secondly, the boundary of holographic codes exhibits a quasiperiodic self-similarity [54, 55], allowing one to rescale the number of physical and logical qubits using local inflation rules that also determine the code’s rate and distance scaling [37, 56], both of which fare better than in constructions of topological quantum codes [57, 58, 59], as well as in many of the recent proposals for quantum low-density parity-check (qLDPC) codes [60]. Thirdly, it is known that most seed tensors for holographic codes can be easily mapped to graph states, implying that efficient preparation schemes likely exist [61, 62, 63, 64]. Lastly, and perhaps most consequentially, it has been shown that holographic codes exhibit high resilience against various noise channels. Indeed, it was first shown in [33] that holographic quantum error correction codes exhibit high thresholds against the quantum erasure channel; subsequently, several works have shown the capability of certain codes constructions to protect against depolarizing and 1-Pauli noise, potentially as well as topological codes [37, 39, 40, 36, 65, 56]. To our knowledge, a systematic biased-noise threshold study for even zero-rate holographic codes has not been performed, let alone for their constant-rate versions111 [41] discusses biased noise in finite-rate HaPPY pentagon codes, but it only captures the non-detectable error probability for a code of fixed length.. As such, our work takes the first step towards understanding holographic codes under more generalized noise channels.
@@
The main ingredient of a holographic quantum code is the seed tensor defining the encoding map for a single tile. Here we consider the zero-rate case with a logical qubit only on the central tile; on the remaining tiles all tensor legs are planar, effectively dispersing the logical information towards the tiling boundary. Note that in generic holographic codes, each tile may hold a logical qubit, leading to a nonzero asymptotic rate as the number $L$ of layers is increased. It was recently proven that holographic codes are an instance of generalized code concatenation [66, 67], and, while copies of the same seed tensor are typically used, one may also employ several different seed tensors in similar spirit to heterogeneous concatenation methods for tree-style concatenated codes [68, 69, 67].
II.2 Tensor-Network Decoding
Contracting a tensor network representing a holographic-code state with potentially thousands of qubits is computationally intractable; as such, we follow the simplification proposed in [36, 37]. In what follows, we briefly review this technique.
In the tensor-network decoder formalism, we represent a stabilizer operator as elements in $\mathbb{Z}_{4}$. As an example, we represent the stabilizer $X_{1}Z_{2}Y_{3}Y_{4}X_{5}I_{6}$ for the $\llbracket 6,1,3\rrbracket$ code as the vector $[1,3,2,2,1,0]$. In a stabilizer code, we naturally have a prescription for defining four logical operators: $\{\mathsf{\bar{I}},\mathsf{\bar{X}},\mathsf{\bar{Y}},\mathsf{\bar{Z}}\}\in\mathsf{L}$. These operators can be used to define logical equivalence classes, permitting us to formulate a rank-$n$ tensor as
$\mathsf{T}(\mathsf{L})_{\alpha_{1},\cdots,\alpha_{n}}:=\begin{cases}1&\text{if$P_{\alpha_{1}}\otimes\cdots\otimes P_{\alpha_{n}}\in\mathsf{SL}$}\\ 0&\text{otherwise}\end{cases}~{}.$ (1)
Here, we take $\alpha_{1},\cdots,\alpha_{n}$ to be the mapping of a given stabilizer or logical operator to $\mathbb{Z}_{4}$, as shown above; $P_{\alpha_{1}}\otimes\cdots\otimes P_{\alpha_{n}}$ represents the Pauli stabilizer formed from the inverse mapping, and $\mathsf{SL}$ is the logical coset of $\mathsf{S}$ with respect to $\mathsf{L}$.
As a side remark, the tensors of Equation 1 are not, strictly speaking, isometric. However, they facilitate efficient contraction, provided that the proper index contraction sequence is indicated; as such, the full decoding process which we will describe below exhibits a runtime complexity of $\mathcal{O}(n^{2.37})$ in the best case [70, 71]. More details on the formation of large tensor networks using the aforementioned scheme can be found in [36, 72].
Tensor-network decoder is a maximum-likelihood (ML) decoder, which is shown to be optimal given the error model. The ML decoder generally works by calculating the most-likely correction needed to return the system to the correct code state, given a syndrome. Here, we define a syndrome as the set $\mathbf{s}=[s_{1}\cdots s_{n-k}]$ of $\mathbb{Z}_{2}$ parity results from measuring each stabilizer. An error $\mathsf{E}$ is initialized as a probability vector corresponding to each individual physical qubit; this vector is then contracted with the holographic tensor network, permitting us to calculate the syndrome $\mathbf{s}=\mathsf{HE}$, where $\mathsf{H}$ is the parity-check matrix of the holographic code.
After finding the syndrome, we deduce the pure error (or destabilizer) $\mathsf{E}_{\text{pure}}$ by taking the Moore-Penrose inverse (or pseudoinverse) $\mathsf{H}_{\text{pseudo}}^{\dagger}$ of the parity-check matrix, resulting in the equation
$\mathsf{E}_{\text{pure}}=\mathsf{H}_{\text{pseudo}}^{\dagger}\mathbf{s}~{}.$ (2)
Once the pure error is obtained, we proceed to calculate the probability that an error with the form of $\mathsf{E}_{\text{pure}}$ has occurred on the $n$ physical qubits; this is done by contracting the tensor network with tensors that parameterizes the probability of errors on each physical qubit on the boundary. For independently distributed errors, even an exact contraction is efficient. From this procedure, we obtain the expression
$\mathbb{P}(\mathsf{L},\mathbf{s}):=\sum_{S\in\mathsf{S}}\text{Prob}[\mathsf{E}_{\text{pure}}|SL]~{},$ (3)
where $\text{Prob}[\mathsf{E}_{\text{pure}}|SL]$ signifies the probability of a pure error acting on the coset of stabilizer $S\in\mathsf{S}$ with respect to logical operator $L\in\mathsf{L}$. In practice, the decoder must find $L$ such that $\mathbb{P}(\mathsf{L},\mathbf{s})$ is maximized, i.e. $\operatorname*{arg\,max}\big{[}\mathbb{P}(\mathsf{L},\mathbf{s})\big{]}$. As an additional condition, the summation over all logical cosets must be equivalent to unity; that is $\sum_{L\in\mathsf{L}}\sum_{S\in\mathsf{S}}\mathbb{P}(\mathsf{L},\mathbf{s})=1$. As before, we refer the reader to [36, 37, 72] for further details on calculating the logical success rate.
@@ III Results
III.1 Noise Model and Setup
The error model tested in this work is principally the code-capacity error model for Pauli noise. Such an error model takes on the form
$\mathsf{E}(\rho)=(1-p)\rho+p(r_{X}X\rho X+r_{Y}Y\rho Y+r_{Z}Z\rho Z)\ ,$ (4)
with the relative error probabilities $r_{X}+r_{Y}+r_{Z}=1$, typically in the regime of $r_{X}=r_{Y}$ with a bias $\eta=\frac{r_{Z}}{r_{X}+r_{Y}}\geq\frac{1}{2}$ (in the case of $Z$-biased noise). When $\eta=\frac{1}{2}$, one recovers the standard depolarizing noise channel. If we move $\eta\mapsto 0$, then we necessarily move towards 2-Pauli noise; that is, noise channels consisting only of two Pauli operators, such as XZ or XY noise.
The seed tensors and their associated logical operators and their stabilizers are shown in Table 1; we have shown all of the seed tensors’ stabilizers and logical operators studied in this work. Here we have used the notation $XX$ to denote $X\otimes X$ for brevity. Additionally, the final entry for each stabilizer generator (with a bar above) denotes the action on the logical index.
Seed Tensor Stabilizers Logical Operators
Tailored $\llbracket 7,1,3\rrbracket$ $XZIZXII\bar{I},IXZIZXI\bar{I},IIXZIZX\bar{I},XIIXZIZ\bar{I},ZXIIXZI\bar{I},IZXIIXZ\bar{I},ZIZXIIX\bar{I}$ $ZZZZZZZ\bar{Z}$, $XXXXXXX\bar{X}$
CD-Steane $XZZIIIX\bar{I}$, $XIZXZII\bar{I}$, $XIIIZZX\bar{I}$, $ZXXIIIZ\bar{I}$,$ZIXZXII\bar{I}$, $ZIIIXXZ\bar{I}$ $XZZXZZX\bar{X}$, $ZXXZXXZ\bar{Z}$
Threshold calculations were performed using the tensor-network decoder from [36, 37]. These decoders and others will be made public in an upcoming software package for holographic quantum error correction codes [73]. However, for the purposes of understanding the current work, we focus on surveying biased-noise resilience for a swath of known holographic codes, and not on a systematic treatment of tensor-network decoding methodology [70].